OpenStudy (anonymous):

@kymber is twice @sheg age. In 3 years, kymber will be 17. How old is sheg?

OpenStudy (anonymous):

She is 7. 2S = K 2S + 3 = 17 -> S = 7

OpenStudy (anonymous):

After 3 yrs Kymber is 17 yrs old so today age of kymber is = 17-3 = 14 yrs sheg's age is half of kymber so sheg's age is 7 yrs

OpenStudy (anonymous):

yep :)

OpenStudy (kymber):

WHAT IS THIS